Wingate Company borrowed $90,000 on January 2, 2019. This amount plus accrued interest of 6% compounded annually will be repaid at the end of 3 years.

What amount will Wingate repay at the end of the third year? (For calculation purposes, use 5 decimal places as displayed in the factor table provided. Round answer to 2 decimal places, e.g. 52.75.)

Amount $
Answer #1

We use the formula:  
A=P(1+r/100)^n
where   
A=future value
P=present value  
r=rate of interest
n=time period.

A=90,000*(1.06)^3

=90,000*1.19102

=$107191.8